OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through the point (4, -1) and is perpendicular to the line whose equation is 2x - y - 7 = 0?

OpenStudy (anonymous):

2x - y - 7 = 0 y = 2x -7 slope of this line = 2 so slope of required line = -1/2 use the general formula y-y1 = m(x-x1) so y - -1 = (-1/2) (x-4) y + 1 = (-1/2)x + 2 y = (-1/2)x + 1